# LocDiffusion: Identifying Locations on Earth by Diffusing in the Hilbert Space

### v1

@@ 5.1 Experimental Setup
We follow the experimental setup of GeoCLIP (Vivanco et al., 2023), a widely used image geolocalization model, for a fair comparison. The training dataset is MP16 (MediaEval Placing Tasks 2016, (Larson et al., 2017)) containing 4.72 million geotagged images. The test datasets is Im2GPS3k (Hays & Efros, 2008). The GWS15k (Clark et al., 2023) reported in GeoCLIP is unfortunately not publicly available. Note that the test sets Im2GPS3k have similar distributions to MP16, and more importantly their data points might overlap with those in MP16, which benefits retrieval-based approach. we first conduct In-Domain Gallery experiments. To compare models without such data distribution artifacts we then conduct generalizability experiments where the anchor points of LocDiffusion and image gallery are sampled from even distributions.
@@
For each test image, our model conditionally generates 16 locations and uses their geographical center as the prediction. Then we count how many predictions fall into the neighborhoods of the ground-truth locations at different scales (1 km, 25 km, 200 km, 750 km and 2500 km) respectively. In Appendix A.3 we present a detailed analysis showing that larger $L$ can improve LocDiffusion resolution. However, we also find that there is a limit of $L=47$ before hitting numerical stability issues (Figure 4).
5.2 In-Domain Gallery Results
We first study the case where training and testing locations come from overlapping distributions. Table 1 summarizes the geolocalization performance of our LocDiffusion model against baselines. On Continent (2500 km), Country (750 km), and Region (200 km) levels, our model can outperform GeoCLIP on the Im2GPS3k dataset. However, on the finer scales (1 km and 25 km), LocDiffusion shows inferior performance to GeoCLIP. Moreover, LocDiffusion cannot outperform PIGEON (Hays & Efros, 2008) on this dataset. We hypothesize that this is due to the limited spatial resolution of SHDD representations. With $L=47$, the intrinsic variance of the SHDD decoder is around 200 km, making predictions on the 1 km and 25 km scales less reliable. To further increase the resolution without increasing $L$, we develop a hybrid approach (LocDiffusion+GeoCLIP in Table 1) that combines the advantages of LocDiffusion and retrieval-based models such as GeoCLIP: we use LocDiffusion to generate candidate locations, and restrict the retrieval of GeoCLIP to the 200 km radius region around the candidate locations. Please see Appendix A.4 for a detailed description of this hybrid model. We can see that LocDiffusion+GeoCLIP has much higher performances across all scales compared to the original LocDiffusion and GeoCLIP on the Im2GPS3k dataset. It can outperform all baselines on Im2GPS3k datasets except the city scale in which it remains competitive with PIGEON (Haas et al., 2024). Nevertheless, we believe leveraging gallery information at smaller scales (the strategy of PIGEON) can lead to even better performance, which we leave as future work.
Table 1: In domain gallery results. The evaluation setup is identical to GeoCLIP (Vivanco et al., 2023). The GeoCLIP model retrieves locations from the 100k gallery provided in its code-base which aligns with the spatial distribution of test images. $L$ is the degree of SHDD representations used in our model. Bold numbers denote the best performance on the corresponding dataset.
@@ 5.2 In-Domain Gallery Results
Dataset Model Street City Region Country Continent
1 km 25 km 200 km 750 km 2500 km
Im2GPS3k [L]kNN, $\sigma$=4 (Vo et al., 2017) 7.2 19.4 26.9 38.9 55.9
PlaNet (Weyand et al., 2016) 8.5 24.8 34.3 48.4 64.6
CPlaNet (Seo et al., 2018) 10.2 26.5 34.6 48.6 64.6
ISNs (Muller-Budack et al., 2018) 10.5 28.0 36.6 49.7 66.0
Translocator (Pramanick et al., 2022b) 11.8 31.1 46.7 58.9 80.1
GeoDecoder (Clark et al., 2023) 12.8 33.5 45.9 61.0 76.1
GeoCLIP (Vivanco et al., 2023) 14.1 34.5 50.7 69.7 83.8
PIGEON (Haas et al., 2024) 11.3 36.7 53.8 72.4 85.3
LocDiffusion ($L$=47) 10.9 34.0 53.3 72.5 85.2
LocDiffusion ($L$=47) + GeoCLIP 14.4 35.8 56.4 73.3 85.5
@@
Table 2: Generalizability experiment. Bold denotes the best results obtained in the given model and gallery/anchor setting. Numbers in the brackets denote relative performance degradation compared to the prior knowledge gallery/anchor.
@@
5.3 Generalizability Experiment Results
The biggest advantage of generative geolocalization over traditional classification/retrieval-based geolocalization methods is that it completely gets rid of predefined spatial classes and location galleries. As is noted in (Vivanco et al., 2023), the performance of retrieval-based geolocalization methods depends heavily on the quality of the gallery – i.e., how well the candidate locations in the gallery cover the test locations. For example, GeoCLIP uses a 100k gallery with locations drawn from MP16 training data. When using this gallery for the GWS15k dataset, the performance drops because there are unseen locations. It was also noticed that GeoCLIP’s performance drops when an evenly sampled grid on Earth is used. At small scales, this is explainable because the grids are too coarse to predict 1 km to 25 km objects. However, at large scales, the performance of GeoCLIP also drops significantly which is unexpected. See the results in Table 2. With 1 million grid points, the average distance between two candidates is less than 30 km. However, the performance of GeoCLIP at the 200 km, 750 km, and 2500 km scales (way larger than 30 km) is still much lower than the performance when using 100K MP16 gallery locations. It indicates that the decline in performance is due to GeoCLIP’s weak generalization to new, unseen locations. We can see that the gallery has a strong inductive bias that narrows the spatial scope, and makes the retrieval model easier to overfit, but hurts its spatial generalizability.
Our LocDiffusion model, though also uses anchor points for decoding (training is random), is almost unaffected by the choice of anchor points. To align with GeoCLIP, we use the same MP16 gallery and evenly sample grid points as decoding anchor points. We can see, at the smaller scales, just like GeoCLIP, introducing the MP16 gallery helps improve the accuracy because its spatial inductive bias helps offset the vagueness of decoding. However, at larger scales, the performance of LocDiffusion is almost independent of the choice of anchors – both the way how we pick the anchor points (MP16 or even grid) and the total number of anchor points (from 21k to 1M). It is a strong indicator of better spatial generalizability for LocDiffusion.
@@ 6 Conclusion, Limitations and Future Work
In this paper, we propose a novel SHDD encoding-decoding framework that enables latent diffusion for spherical location generation. We also propose a CS-UNet architecture to learn conditional diffusion and train a LocDiffusion model that addresses the image geolocalization task via generation. It achieves competitive geolocalization performance and demonstrates significantly better spatial generalizability.
One limitation is that to accurately generate locations at finer scales, we need to quadratically increase the SHDD encoding dimension, which is computationally demanding. We show that by combining LocDiffusion and a retrieval-based image geolocalization method can significantly improve model performances across all scales. In the future, we aim to leverage gallery during the diffusion process and explore solutions like hierarchical generation that reduce SHDD space complexity to linear.
Impact Statement
This work introduces LocDiffusion, a novel diffusion-based geolocalization method that predicts locations on a spherical surface with high spatial resolution, overcoming limitations of grid-based and gallery-based approaches that restrict spatial accuracy. By leveraging the proposed SHDD representation and a latent diffusion model, LocDiffusion enables precise geolocation predictions for tasks such as wildlife tracking, disaster response, and processing remote sensing images. This research advances geospatial machine learning by providing a scalable and accurate solution for location prediction, with potential to improve environmental monitoring and urban planning.
@@ Appendix A Appendix
A.1 Sparsity of Existing Positional Encoding Methods
Almost all location encoders can be formulated as the following equation (Wu et al., 2024):
[…]
1) $tile$ is a vanilla location encoder used by many pioneering studies(Berg et al., 2014; Adams et al., 2015; Tang et al., 2015). It divides geographic regions into discrete global grids based on longitude and latitude and learns corresponding partition embeddings based on the grid cell indicator vectors.
[…]
8) $NeRF$ can be viewed as a multiscale version of $xyz$ by employing Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) as its position encoder.
9) Sphere2Vec (Mai et al., 2023b), including Sphere2Vec-$sphereC$, Sphere2Vec-$sphereC+$, Sphere2Vec-$sphereM$, Sphere2Vec-$sphereM+$, and Sphere2Vec-$dfs$, is a series of multi-scale location encoders for spherical surface based on Double Fourier Sphere (DFS) and Space2Vec. The multi-scale representation of Sphere2Vec is achieved by one-to-one mapping from each point $x_{i}=(\lambda_{i},\varphi_{i})\in\mathbb{S}^{2}$ with $\mathit{S}$ be the total number of scales. They are the first location encoder series that preserves the spherical surface distance between any two points to our knowledge.
10) Siren (SH) (Rußwurm et al., 2024) is a more recently proposed spherical location encoder, which claims a learned Double Fourier Sphere location encoder. It uses spherical harmonic basis functions as the position encoder $\mathbb{PE}^{\textit{Siren (SH)}}(\vec{x})$, followed by a sinusoidal representation network (SirenNets) as the $\mathbf{NN}()$.

### v2

@@ 5.1 Main Results
We first follow the experimental setup of GeoCLIP geoclip , a widely used image geolocalization model, for a fair comparison. The training dataset is MP16 (MediaEval Placing Tasks 2016 larson2017benchmarking ) containing 4.72 million geotagged images. The test datasets are 3 global-scale image geolocalization datasets – Im2GPS3k (hays2008im2gps, ), YFCC26k (thomee2016yfcc100m, ), and GWS15k (clark2023we, ). Note that the test datasets Im2GPS3k and YFCC-26k have similar distributions to MP16, and more importantly, their data points might overlap with those in MP16, which benefits retrieval-based approaches geoclip .
@@
During model inference, we run multiple (16) samplings given different augmentations of each test image and use their geographical center as the prediction. This is a successful strategy proven by GeoCLIP(geoclip, ) and SimCLR(chen2020simple, ). For example, in our experiments, without this augmentation and averaging step, the 1km accuracy of GeoCLIP on Im2GPS3K may drop from 14% to below 10%. Then we count how many predictions fall into the neighborhoods of the ground-truth locations at different scales respectively, which are denoted as Street (1 km), City (25 km), Region (200 km), County (750 km), and Continent (2500 km). See Appendix A.4, for a detailed ablation study on $L$. Larger $L$ can improve LocDiff resolution, but there is a limit of $L=47$ before hitting numerical stability issues (Figure 5). We leave numerical stability issue to our future work.
Table 1: Main evaluation results with the GeoCLIP geoclip eval setup. $L$ is the degree of SHDD representations used in our model. LocDiff-H stands for the hybrid approach of LocDiff plus GeoClip. “Coun.” and “Cont.” indicate Country and Continent scales. Bold and underline numbers denote the best and second best performance on the corresponding dataset. - means not reported.
Model Im2GPS3k YFCC-26k GWS15k
Street City Region Coun. Cont. Street City Region Coun. Cont. Street City Region Coun. Cont.
1 km 25 km 200 km 750 km 2.5k km 1 km 25 km 200 km 750 km 2.5k km 1 km 25 km 200 km 750 km 2.5k km
[L]kNN, $\sigma$=4 vo2017revisiting 7.2 19.4 26.9 38.9 55.9 - - - - - - - - - -
PlaNet weyand2016planet 8.5 24.8 34.3 48.4 64.6 4.4 11.0 16.9 28.5 47.7 - - - - -
CPlaNet seo2018cplanet 10.2 26.5 34.6 48.6 64.6 - - - - - - - - - -
ISNs muller2018geolocation 10.5 28.0 36.6 49.7 66.0 5.3 12.3 19.0 31.9 50.7 0.05 0.6 4.2 15.5 38.5
Translocator pramanick2022world 11.8 31.1 46.7 58.9 80.1 7.2 17.8 28.0 41.3 60.6 0.5 1.1 8.0 25.5 48.3
GeoDecoder clark2023we 12.8 33.5 45.9 61.0 76.1 10.1 23.9 34.1 49.6 69 0.7 1.5 8.7 26.9 50.5
GeoCLIP geoclip 14.1 34.5 50.7 69.7 83.8 11.6 22.2 36.7 57.5 76.0 0.6 3.1 16.9 45.7 74.1
PIGEON Haas_2024_CVPR 11.3 36.7 53.8 72.4 85.3 10.5 25.8 42.7 63.2 79.0 0.7 9.2 31.2 65.7 85.1
LocDiff (LL=47) 10.9 34.0 53.3 72.5 85.2 9.6 22.8 37.5 58.6 76.8 2.1 12.4 33.7 67.0 85.0
LocDiff-H (LL=23) 15.3 36.5 56.4 75.2 87.4 13.2 26.0 41.9 64.5 80.3 0.9 7.4 33.5 66.2 85.0
Table 1 summarizes the performance of different models on these three datasets. In addition to the pure LocDiff model, to further increase the resolution without increasing $L$, we develop a hybrid approach denoted as LocDiff-H in Table 1 that combines the advantages of LocDiff and retrieval-based models such as GeoCLIP – we use LocDiff to generate candidate locations, and restrict the retrieval of GeoCLIP to the 200 km radius region around the candidate locations. Please see Appendix A.6 for a detailed description of this hybrid model. From Table 1, we can see that LocDiff-H can outperform all baselines and LocDiff across almost all spatial scales on both Im2GPS3k and YFCC-26k datasets while remaining competitive for the city scale metric on Im2GPS3k and region scale metric on YFCC-26k. On the GWS15k dataset, LocDiff shows the best performance while LocDiff-H underperforms LocDiff, especially on higher spatial scale metrics (e.g., street and city scales). We hypothesize that this is because, unlike Im2GPS3k and YFCC-26k, the spatial distribution of GWS15k is very different from that of the MP16 training dataset. The spatial inductive bias brought by GeoCLIP will not benefit but hurt the performance of LocDiff-H on GWS15k, especially for fine-scale metrics.
Table 2: Comparison with existing generative methods. The evaluation setup is identical to Dufour et. al dufour2024world80timestepsgenerative . $L$ is the degree of SHDD representations used in our model. Bold numbers denote the best performance on the corresponding dataset. - means not reported.
@@
Dataset Model City Region Country Continent Density Coverage
25 km 200 km 750 km 2500 km
OSV-5M vMF dufour2024world80timestepsgenerative 0.6 17.2 52.7 - - -
vMFMix izbicki2020exploiting 0.3 11.1 34.2 - - -
Diff $\mathbb{R}^{3}$ dufour2024world80timestepsgenerative 3.6 40.9 75.9 - 0.752 0.568
FM $\mathbb{R}^{3}$ dufour2024world80timestepsgenerative 4.2 40.0 74.9 - 0.799 0.575
RFM $\mathcal{S}_{2}$ dufour2024world80timestepsgenerative 5.4 44.2 76.2 - 0.797 0.590
LocDiff (LL=47) 11.0 46.3 77.0 88.2 0.795 0.572
YFCC-4k vMF dufour2024world80timestepsgenerative 4.8 15.0 30.9 53.4 - -
vMFMix izbicki2020exploiting 0.4 8.8 20.9 41.0 - -
Diff $\mathbb{R}^{3}$ dufour2024world80timestepsgenerative 11.1 37.7 54.7 71.9 0.959 0.837
FM $\mathbb{R}^{3}$ dufour2024world80timestepsgenerative 22.1 35.0 53.2 73.1 1.037 0.920
RFM $\mathcal{S}_{2}$ dufour2024world80timestepsgenerative 23.7 36.4 54.5 73.6 1.060 0.926
RFM $\mathcal{S}_{2}$ (10M) dufour2024world80timestepsgenerative 33.5 45.3 61.1 77.7 - -
LocDiff (LL=47) 33.3 46.7 65.2 79.9 1.072 0.915
Several generative models have been developed recently for image geolocalization including Diff $\mathbb{R}^{3}$ dufour2024world80timestepsgenerative , FM $\mathbb{R}^{3}$ dufour2024world80timestepsgenerative , and RFM $\mathcal{S}_{2}$ dufour2024world80timestepsgenerative . Since these models were evaluated in different datasets and experimental setups, in order to conduct a fair comparison, we follow the evaluation setup in dufour2024world80timestepsgenerative and compare our LocDiff against these generative geolocalization models on two datasets – OSV-5M astruc2024osv5m and YFCC-4k vo2017revisiting . The results are shown in Table 2. We can see that LocDiff can outperform all generative geolocalization models on both datasets which clearly showcases the advantages of our multi-scale latent diffusion approach compared with those diffusion models directly applied on the original coordinate space dufour2024world80timestepsgenerative .
5.2 Generalizability Experiment Results
The key advantage of generative geolocalization over traditional classification/retrieval-based geolocalization methods is that it completely gets rid of predefined spatial classes and location galleries. As noted in geoclip , the performance of retrieval-based geolocalization methods depends heavily on the quality of the gallery – i.e., how well the candidate locations in the gallery cover the test locations. For example, GeoCLIP uses a 100k gallery with locations drawn from MP16 training data. When using this gallery for the GWS15k dataset, the performance drops due the spatial distribution mismatch between MP16 and GWS15k. As shown in Table 3, GeoCLIP’s performance significantly drops when an evenly sampled grid on Earth is used instead of the default image gallery. At small scales, this is explainable because the grids are too coarse to differentiate 1 km to 25 km objects. However, at large scales, the performance of GeoCLIP also drops significantly which is unexpected. With 1 million grid points, the average distance between two candidates is less than 30 km. However, the performance of GeoCLIP at the 200 km, 750 km, and 2500 km scales (way larger than 30 km) is still much lower than the performance when using 100K MP16 gallery locations. It indicates that the decline in performance is due to GeoCLIP’s weak generalization to new, unseen locations. We can see that the gallery has a strong inductive bias that narrows the spatial scope, and makes the retrieval model easier to overfit, but hurts its spatial generalizability.
Our LocDiff model, while also using anchor points for decoding (training is random), is almost unaffected by the choice of anchor points. To align with GeoCLIP, we compare the same MP16 gallery and evenly sample grid points as decoding anchor points. We can see, at the smaller scales, just like GeoCLIP, introducing the MP16 gallery helps improve the accuracy because its spatial inductive bias helps offset the vagueness of decoding. However, at larger scales, the performance of LocDiff is almost independent of the choice of anchors – both the way how we pick the anchor points (MP16 or even grid) and the total number of anchor points (from 21k to 1M). It is a strong indicator of better spatial generalizability for LocDiff.
@@ 5.2 Generalizability Experiment Results
Table 3: Generalizability experiment. Numbers outside of the brackets denote geolocalization accuracies. Numbers in the brackets denote relative performance degradation compared to the prior knowledge gallery/anchor. Bold numbers denote the largest drop in performance with the given gallery/anchor setting.
@@
5.3 Computational Efficiency
Though the mathematical derivation of SHDD encoding/decoding seems complicated, our method does NOT introduce much computational cost. First, the SHDD encoding/decoding operations are both deterministic, closed-form. For example, the $d$-dimensional ($d=256,576,1024$) SHDD encoding of a point is just a parallel evaluation of a list of two-variable functions, whose time complexity is $O(1)$ and space complexity is $O(d)$. Second, during training, the SHDD encodings can be precomputed and used as an embedding lookup table and there are no further computational overheads. SHDD decoding is implemented as a matrix multiplication followed by an argmax operation, which is also very efficient. See Table 4.
Encoding DDPM Decoding
CPU 0.0003s per location GPU 0.056s per image (200 steps, 16 augmentations) GPU 0.0012s per batch (512)
Table 4: Time efficiency for unit operations in LocDiff
We wish to emphasize that using an SHDD encoded multi-scale representation helps the diffusion process to converge faster, because each dimension encodes information at the respective spatial scales. For example, if at a certain step, the generated SHDD representation is 500km within the ground-truth, then the dimensions with larger than 500km scales will remain mostly intact, and only the finer dimensions need to be altered. In practice, our model takes about 2 million steps to converge on YFCC, while the best results reported in dufour2024world80timestepsgenerative take 10 million steps.
@@ 6 Conclusion, Limitations and Future Work
In this paper, we propose a novel SHDD encoding-decoding framework that enables multi-scale latent diffusion for spherical location generation. We also propose a CS-UNet architecture to learn conditional diffusion and train a LocDiff model that addresses the image geolocalization task via location generation. It achieves the state-of-the-art geolocalization performance over all existing baselines on 5 global-scale image geolocalization datasets and demonstrates significantly better spatial generalizability. In the future, we aim to leverage the gallery during the diffusion process and explore different ways to further speed up decoding.
Acknowledgements
This work is mainly funded by the National Science Foundation under Grant No. # 2521631 – A Statistics-Based Geographic Bias Quantification and Debiasing Framework for GeoAI and Foundation Models. Gengchen Mai acknowledges support from the Microsoft Research Accelerate Foundation Models Academic Research (AFMR) Initiative and University of Texas at Austin Library Map & Geospatial Collections Explorer Fellowship. Zeping Liu acknowledges the support of the Amazon AI PhD fellowship. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
@@ Appendix A Appendix
A.1 Example of Diffusion Steps
Figure 3: Illustration of how the spherical probability mass concentration (mapped to a plane) corresponding to the SHDD encodings changes along the backward process at step=0, 10, 20, 100, 200, respectively. The more bright, the more probability mass.
A.2 Sparsity of Existing Positional Encoding Methods
Almost all location encoders can be formulated as the following equation [48]:
$\mathit{Enc}(\theta,\phi)=\mathbf{NN}(\mathbb{PE}(\theta,\phi)),$ (15)
$\mathbb{PE}()$ is a position encoder that transforms the location $\mathbf{p}=(\theta,\phi)$ into a $\mathit{W}$-dimensional vector, referred to as the position embedding. The neural network $\mathbf{NN}():\mathbb{R}^{W}\rightarrow\mathbb{R}^{d}$ is a learnable function that maps the position embedding $\mathbb{PE}(\theta,\phi)\in\mathbb{R}^{W}$ to the location embedding $\mathit{Enc}(\theta,\phi)\in\mathbb{R}^{d}$.
1) tiletile is a vanilla location encoder used by many pioneering studies[4, 1, 43]. It divides geographic regions into discrete global grids based on longitude and latitude and learns corresponding partition embeddings based on the grid cell indicator vectors.
2) wrapwrap [22] is a sinusoidal location encoder, normalizing latitude and longitude and processing with sinusoidal functions before feeding into $\mathbf{NN}^{\mathit{wrap}}()$, which is composed of four residual blocks implemented through linear layers.
3) $wrap+ffn$ [27] is a variant of wrapwrap that substitutes $\mathbf{NN}^{\mathit{wrap}}()$ with $\mathbf{NN}^{\mathit{ffn}}()$, a simple FFN.
4) rbfrbf [25] is a kernel-based location encoder. It randomly selects $W$ points from the training dataset as Radial Basis Function (RBF) anchor points. It then applies Gaussian kernels to each anchor points.Each input point $\vec{x}_{i}$ is represented as a $W$-dimension feature vector using these kernels, which is then processed by $\mathbf{NN}^{\mathit{ffn}}()$.
5) rffrff stands for Random Fourier Features [34] and it is another kernel-based location encoder. It first encodes location $\vec{x}$ into a $W$ dimension vector - $\mathbb{PE}^{rff}(\vec{x})=\varphi(\vec{x})$. Each component of $\varphi(\vec{x})$ first projects $\vec{x}$ into a random direction $\omega_{i}$ and makes a shift by $b_{i}$. Then it wraps this line onto a unit circle in $\mathbb{R}^{2}$ with the cosine function. $\mathbb{PE}^{rff}(\vec{x})$ is further fed into $\mathbf{NN}^{\mathit{ffn}}()$ to get a location embedding.
6) Space2Vec-gridgrid and Space2Vec-theorytheory [25] are two versions of sinusoidal multi-scale location encoders on 2D Euclidean space. Both of them implement the position encoder $\mathbb{PE}(\vec{x})$ as performing a Fourier transformation on a 2D Euclidean space then fed into the $\mathbf{NN}^{\mathit{ffn}}()$. Space2Vec-gridgrid treats $x=(\lambda,\varphi)$ as a 2D coordinate while Space2Vec-theorytheory be simulated by summing three cosine grating functions oriented 60 degree apart.
7) xyzxyz [27] is a vanilla 3D location encoder, converting the lat-lon spherical coordinates into 3D Cartesian coordinates centered at the sphere center with position encoder $\mathbb{PE}^{xyz}(\vec{x})$, then feeds the 3D coordinates into an MLP $\mathbf{NN}^{\mathit{ffn}}()$.
8) $NeRF$ can be viewed as a multiscale version of xyzxyz by employing Neural Radiance Fields (NeRF) [28] as its position encoder.
9) Sphere2Vec [27], including Sphere2Vec-$sphereC$, Sphere2Vec-$sphereC+$, Sphere2Vec-sphereMsphereM, Sphere2Vec-$sphereM+$, and Sphere2Vec-dfsdfs, is a series of multi-scale location encoders for spherical surface based on Double Fourier Sphere (DFS) and Space2Vec. The multi-scale representation of Sphere2Vec is achieved by one-to-one mapping from each point $x_{i}=(\lambda_{i},\varphi_{i})\in\mathbb{S}^{2}$ with $\mathit{S}$ be the total number of scales. They are the first location encoder series that preserves the spherical surface distance between any two points to our knowledge.
10) Siren (SH) [37] is a more recently proposed spherical location encoder, which claims a learned Double Fourier Sphere location encoder. It uses spherical harmonic basis functions as the position encoder $\mathbb{PE}^{\textit{Siren (SH)}}(\vec{x})$, followed by a sinusoidal representation network (SirenNets) as the $\mathbf{NN}()$.